System of equations:
-7x+3y=-65
-7x+10y=-135
Write an equation that results from subtracting two equations

Answer by ikleyn(52756)

.
System of equations:
-7x+3y=-65
-7x+10y=-135
Write an equation that results from subtracting two equations
~~~~~~~~~~~~~~~~~~~~~~~~~

If you subtract first equation from the second equation, the resulting equation will be


    10y - 3y = -135 - (-65),

or

       7y    = -135 + 65,

or

       7y    =    -70       <<<---===   final

Subtracted and explained.